Find the vector equation of the plane determined by the points (0,2,1),(2,1,0),(1,-1,0). Further, check whether the line r=(1+2α)I + (2-3α)j - (1+5α)k intersects this plane. If it intersects, find the point of intersection. If the line and plane do not intersects,find the equation of another line that intersects this plane.


1
Expert's answer
2021-05-25T08:59:28-0400

"A(0,2,1), B(2,1,0),C(1,-1,0)\\
\begin{vmatrix}
 x-0 & y-2&z-1 \\
 2-0 & 1-2&0-1\\
1-0&-1-2&0-1
\end{vmatrix}=0\\"

"\begin{vmatrix}
 x & y-2&z-1 \\
 2 & -1&-1\\
1&-3&-1
\end{vmatrix}=0\\"

"-2x+y-5z+3=0"

The vector equation of the plane:

"(0,2,1)+v(2-0,1-2,0-1)+\\
+u(1-0,-1-2,0-1)=\\
=(2v+u,2-v-3u,1-v-u)"

"r=(1+2a)i+(2-3a)j-(1+5a)k=\\
=(1+2a,2-3a,-(1+5a))\\
-2(1+2a)+(2-3a)-5(-1-5a)+3=0\\
18a+8=0\\
a=-\frac{4}{9}\\
x=1+2\cdot\frac{-4}{9}=\frac{1}{9}\\
y=2-3\cdot \frac{-4}{9}=\frac{10}{3}\\
z=-(1+5\cdot \frac{-4}{9})=\frac{11}{9}\\
(\frac{1}{9}, \frac{10}{3},\frac{11}{9})"